# Adapting Large Language Models for Education: Foundational Capabilities, Potentials, and Challenges

### v2

@@ 2.1.2 Challenging Mathematics Problems
Despite occasional errors in basic mathematical operations, the expectation for LLMs to solve more complex problems remains high, and the LLMs’ capabilities in this area continue to develop. The ability to handle college-level mathematics is particularly beneficial for senior students’ learning, offering significant assistance in understanding challenging concepts. In recent years, more and more work has been proposed to develop this ability (Luo et al., 2023a; Yu et al., 2023; Lewkowycz et al., 2022; Wang et al., 2023a). Wang et al. (2023a) introduced a benchmark SCIBENCH, which contains collegiate-level scientific problems from mathematics, chemistry, and physics textbooks. Luo et al. (2023a) tried to enhance the mathematical reasoning abilities of Llama-2 by applying Reinforcement Learning from Evol-Instruct Feedback (RLEIF). Zhou et al. (2023) tried to enhance GPT-4’s mathematical ability by encouraging it to use code to self-verify its answers. This approach leads to a significant improvement in the zero-shot accuracy of mathematical problem-solving.
2.1.3 Problems Involving Multi-Modal Information
@@
Multi-modal inputs like geometric shapes in math problems require LLMs to understand text and image information for solutions. Research in multi-modal LLMs for mathematical reasoning is emerging (Lu et al., 2023; Chen et al., 2022b; Peng et al., 2023b). This type of task poses a high requirement for the training data’s formation and quality. Chen et al. (2022b) introduced a Unified Geometry problem benchmark combining calculation and proving tasks. Based on this dataset, the study presented a framework capable of simultaneously solving calculations and proving tasks through a sequence generation approach. Lu et al. (2023) proposed MATHVISTA, a benchmark for diverse mathematical and visual challenges. Frameworks for this task usually contain separate encoders for different modalities. For geometric problems, Zhang et al. (2023b) converted diagrams into text clauses, using a convolutional neural network and a language model for encoding and a GRU-based framework for answer generation. Additionally, Lu et al. (2022a) presented the Tabular Math Word Problems (TABWP) dataset, requiring textual and tabular data reasoning, and introduced PROMPTPG, a policy gradient-based selector for training and prompt construction for test samples.
2.1.4 Mathematical Proof
@@
Unlike other types of mathematical problems where LLMs primarily focu on reasoning and providing answers, LLMs’ role in mathematical proofs emphasizes the integration with proof assistants such as Coq (Barras et al., 1997), Isabelle (Nipkow et al., 2002), and Lean (de Moura et al., 2015). These proof assistants correspond to specific programming languages, requiring users to input proofs in these languages, after which the assistant can verify the proof’s correctness.
@@
Many LLM-based methods are proposed to help theorem proving (Jiang et al., 2023; Liu et al., 2023b; Xiong et al., 2023). Based on these proof assistants, there are two main approaches to utilizing LLMs for mathematical proofs. The first approach is formal proof search, exemplified by models like GPT-f (Polu and Sutskever, 2020), which involves prompting LLMs to produce the next proof step(also called ’tactic’ in proof assistants) based on the current proof state and some optional context. Following, Thor (Jiang et al., 2022b) was further proposed to help select the premise for theorem proving. Yang et al. (2023b) introduced an open-source framework named LeanDojo based on the Lean proof assistant. This framework comprises data, toolkits, models, and benchmarks, and it has led to the development of ReProver (Retrieval-Augmented Prover), which enhances proof accuracy by using retrieval methods to extract premises for LLMs to base their mathematical proofs.
@@
The second is natural proof translation, also known as autoformalization, represented by models like DSP (Jiang et al., 2022a), which converts math proofs written in natural language into formalized versions. This approach primarily addresses the issue of insufficient data for formal mathematical proofs. By utilizing autoformalization, a significant increase in this type of data can be achieved. Consequently, this approach enhances the proof-generating capabilities of neural provers that have been fine-tuned on this expanded dataset. The strategy leverages the transformation of existing proofs into a formal structure, thereby enriching the training material for machine learning models in the field of mathematical proof generation. For example, Wu et al. (2022) employed LLMs for autoformalization, transforming mathematical proofs and problems expressed in natural language into formal specifications and proofs in Isabelle language. The generated data was used to train a neural theorem prover, enhancing the effectiveness of the original prover. Follow it, Cunningham et al. (2023) utilized an encoder-decoder framework based on the universal transformer architecture, converting both problem statements and mathematical proofs written in LaTeX into the language of the Coq interactive prover.
@@
2.2 Writing
@@ 2.2.1 Text Summarization
Text summarization, which condenses long texts into summaries, challenges LLMs to understand diverse content like news and cross-lingual texts. It is evident that traditional fine-tuning methods are less effective with the advent of more advanced LLMs (Pu et al., 2023; Liu et al., 2022). Pu et al. (2023) and Liang et al. (2022) showed LLMs like ChatGPT initially lag behind fine-tuned models like T5 (Raffel et al., 2020) and BART (Lewis et al., 2019) in ROUGE scores (Lin, 2004) for text summarization. However, when human judges evaluate overall quality, LLMs outperform fine-tuned models and even standard human summaries, superior in aspects like factual consistency, fluency, and diversity. This discovery underscores the limitations of traditional evaluation methods and suggests a need for new paradigms to guide summarization tasks in the LLM era. For example, BRIO (Liu et al., 2022) implemented a ranking task to foster more diverse summarizations. Furthermore, Liu et al. (2023e) utilized a GPT model based on BRIO to directly generate training data to guide the learning process of other models, which is similar to the process of RLHF (Stiennon et al., 2020).
2.2.2 Grammatical Error Correction
@@
We are well aware of the remarkable capability of LLMs to generate fluent and coherent conversations. However, from an educational perspective, the importance of producing grammatically correct dialogues cannot be overstated, especially for students learning a new language. The correctness of grammar in conversations plays a vital role in language acquisition, providing students with reliable examples to emulate and learn from. Numerous studies have evaluated the effectiveness of LLMs in grammatical error correction. Several works (Wu et al., 2023; Fang et al., 2023; Mohammed et al., 2023) first evaluated the error correction performance of closed-source LLMs such as ChatGPT. Although there exists a pronounced gap between ChatGPT and the previous SOTA (Omelianchuk et al., 2020; Grammarly, 2023) models on the overall F0.5 metric, closer analysis shows that ChatGPT underperforms other models in terms of precision but far exceeds other models in terms of recall. That said, LLMs like ChatGPT are good at error detection. A detailed manual analysis of ChatGPT’s outputs revealed that, in most cases, it maintained grammar accuracy better than the previous methods. However, it often overcorrects sentences to increase diversity and fluency, resulting in a decrease in precision. Nevertheless, it demonstrates that LLMs are adept at generating diverse and accurate content. Some studies have attempted to explore open-source LLMs, such as CoEdit (Raheja et al., 2023), which covered multiple text editing tasks (including GEC) by fine-tuning LLMs to integrate the capabilities brought by these tasks. GrammarGPT (Fan et al., 2023), on the other hand, asked ChatGPT to generate grammatically incorrect sentences to guide the training of other models. Collectively, these studies underscore the prowess of LLMs in generating correct text.
@@ 2.3 Programming
Programming is a process of writing code and correcting code if unexpected results are obtained. Incorporating LLMs in programming education is reshaping the future of AI-assisted programming learning. LLMs could play multiple roles: as instructors providing guidance, as teaching assistants offering personalized tutorials, and as collaborative coding partners. Studies like (Ma et al., 2023a) demonstrated improved performance (17%) and efficiency (13%) among programming novices using LLM-based assistants. Research from (Phung et al., 2023) focused on programming education tasks and benchmarks like (Fu et al., 2023) and (Ding et al., 2023), which were used to evaluate the effectiveness of LLMs. This section mainly discusses the LLMs’ coding development from two perspectives: code writing and code refinement, corresponding to the two stages in programming.
2.3.1 Code Writing
@@
Unlike natural language tasks, generating code requires a more rigorous token syntax and places higher demands on the training stage. A common method to improve LLMs’ performance in generating code is to train or fine-tune them on extensive code datasets (Chen et al., 2021a; Nijkamp et al., 2023). WizardCoder (Luo et al., 2023b) introduced the Evol-Instruct (Xu et al., 2023) method to generate more complex and diverse instruction datasets of code-related tasks. To emulate the iterative process of humans repeatedly modifying and reviewing code, InCoder (Fried et al., 2023) utilized bidirectional encoding instead of left-to-right encoding. In addition to next-token prediction, training or fine-tuning code-aimed LLMs on additional code-related tasks could enhance their programming capabilities. LLMs first learn language patterns and representations from a large amount of text data through unsupervised learning. Then, they could be fine-tuned on labeled code tasks, allowing them to learn more targeted code representations and gain a deeper understanding of code structure and semantics based on the provided labels. CodeT5+ (Wang et al., 2023b) introduced the concepts of unimodal and bimodal alignment, increasing the model’s adaptability to function in different modes for various downstream tasks. During the bimodal alignment phase, the model synchronizes the representations of text-code pairs using multiple tasks, which improves its ability to understand and generate content across different modalities. CodeLlama (Rozière et al., 2023) also applied the multi-task objectives, including autoregression and causal infilling prediction, which achieves better performance among open models. MFTCoder (Liu et al., 2023a) utilized the Multi-Task Learning (MTL) technique and incorporated a training loss computation algorithm to alleviate the instability and imbalance of multi-task training.
2.3.2 Code Refinement

### v3

@@ 3.1.2. Challenging Mathematics Problems
Despite occasional errors in basic mathematical operations, the expectation for LLMs to solve more complex problems remains high, and the LLMs’ capabilities in this area continue to develop. For education, the ability to handle college-level mathematics is particularly beneficial for senior students’ learning, offering significant assistance in understanding challenging concepts. For simple arithmetic problems, LLMs make mistakes mainly due to the gap between text generation and digit calculation. For complex mathematical problems, the challenge arises for requiring LLMs’ advanced symbolic reasoning abilities and domain knowledge. In this regard, LLMs still need further development. Wang et al. (2023a) introduced a benchmark SCIBENCH, which contains collegiate-level scientific problems from mathematics, chemistry, and physics textbooks, while GPT-4 could only get averagely 53.24% on the math part. Furthermore,  Sawada et al. (2023) collected a harder dataset ARB that contains math problems from Harvard PhD comprehensive exams in mathematics, where GPT-4 could only get less than 10% right. All these results demonstrate that LLMs have a lot of room for improvement.
In recent years, more and more work has been proposed to enhance this ability (Luo et al., 2023a; Yu et al., 2023; Lewkowycz et al., 2022; Wang et al., 2023a). Luo et al. (2023a) tried to increase the mathematical reasoning abilities of Llama-2 by applying Reinforcement Learning from Evol-Instruct Feedback (RLEIF) on complex mathematics datasets. The Evol-Instruct method made LLMs to generate easier and harder questions from the original questions to make the LLMs think deeper. In addition to training on high-quality datasets, there are also many efforts that attempt to leverage programming as an external tool to assist in solving mathematical problems. Zhou et al. (2023a) tried to enhance GPT-4’s mathematical ability by encouraging it to use code to self-verify its answers. This approach leads to a significant improvement in the zero-shot accuracy of mathematical problem-solving. ToRA (Gou et al., 2023) divided the process of LLMs solving math problems into a cyclical rationale-action process, where the action involves invoking external tools, including computation libraries and symbolic solvers, thereby amalgamating the analytical prowess of language with the computational efficiency of tools. Overall, solving complex mathematical problems requires reasoning abilities, computational power, and knowledge in the mathematical domain. Fine-tuning on relevant datasets primarily aims to enhance its reasoning capabilities and knowledge on related issues, while computational abilities can be supplemented by invoking external tools, such as calculators or code compilers.
3.1.3. Problems Involving Multi-Modal Information
@@
Multi-modal inputs are common in mathematics problems like geometric questions. They require LLMs to understand text and image information for solutions. Research in multi-modal LLMs for mathematical reasoning is emerging (Lu et al., 2023; Chen et al., 2022a; Peng et al., 2023a). This type of task poses a high requirement for the formation and quality of the training data. Chen et al. (2022a) introduced a Unified Geometry problem benchmark combining calculation and proving tasks. Based on this dataset, the study presented a framework capable of simultaneously solving calculations and proving tasks through a sequence generation approach. Furthermore, Lu et al. (2023) proposed MATHVISTA, a benchmark for diverse mathematical and visual challenges. Zhang et al. (2024) developed a benchmark GeoEval for testing geometry problem-solving ability of LLMs. Their results showed that WizardMath and GPT-4V excels in handling multi-modal mathematics problems.
Various methods have been proposed for this task. For geometric problems, Zhang et al. (2023a) converted diagrams into text clauses, using a convolutional neural network and a language model for encoding and a GRU-based framework for answer generation. Gao et al. (2023) argued that the reason current models fail on solving geometry problems is that they struggle to accurately comprehending basic geometric elements and their relationships. So they built a augmented dataset Geo170K containing high-quality descriptions of geometric information and developed a model G-LLaVA on it, which demonstrated exceptional performance in solving geometric problems, significantly outperforming GPT-4-V on the MathVista benchmark with only 7B parameters.
Besides geometry problems that involves processing images and text, Lu et al. (2022a) presented the Tabular Math Word Problems (TABWP) dataset, requiring textual and tabular data reasoning, and introduced PROMPTPG, a policy gradient-based selector for training and prompt construction for test samples.
3.1.4. Mathematical Proof
@@
Unlike other types of mathematical problems where LLMs primarily focus on providing answers, LLMs’ role in mathematical proofs emphasizes the integration with proof assistants such as Coq (Barras et al., 1997), Isabelle (Nipkow et al., 2002), and Lean (de Moura et al., 2015). These proof assistants correspond to specific programming languages, requiring users to formulate proofs in the required languages, after which the assistant can verify the proof’s correctness. Many LLM-based methods are proposed to help theorem proving (Jiang et al., 2023; Liu et al., 2023e; Xiong et al., 2023). Based on these proof assistants, there are two main approaches to utilizing LLMs for mathematical proofs.
@@
The first approach is formal proof search, exemplified by models like GPT-f (Polu and Sutskever, 2020), which involves prompting LLMs to produce the next proof step (also called ’tactic’ in proof assistants) based on the current proof state and some optional context. In conjunction with proof assistants, it transforms the proof of mathematical propositions into a process of executing actions. Here, an action can be the application of a mathematical theorem or a method of variable substitution, which can transform and decompose the original proposition. The LLM is responsible for generating actions, sampling multiple actions in each round and iterating over multiple rounds to produce a tree structure. It utilizes the proof assistant’s functionality to verify the validity of proofs to evaluate branches, thereby employing tree search methods to find proof strategies. Following GPT-f, Thor (Jiang et al., 2022b) was further proposed to help select the premise for theorem proving. Yang et al. (2023b) introduced an open-source framework named LeanDojo based on the Lean proof assistant. This framework comprises data, toolkits, models, and benchmarks, and it has led to the development of ReProver (Retrieval-Augmented Prover), which enhanced proof accuracy by using retrieval methods to extract premises for LLMs to base their mathematical proofs.
@@
The second is natural proof translation, also known as autoformalization, which is to convert math proofs written in natural language into formalized versions. In such schemes, the responsibility of LLMs is not to generate proof steps. Due to the extremely low prevalence of these proof assistants in the human corpus, LLMs struggle to directly undertake the task of generating proof steps. This approach primarily tackles the challenge of insufficient data for formal mathematical proofs. By leveraging autoformalization, a significant increase in this type of data can be achieved, consequently enhancing the proof-generating capabilities of neural provers that have been fine-tuned on this expanded dataset. Initially,  Wu et al. (2022) demonstrated that LLMs perform well in autoformalization. They employed LLMs for autoformalization, transforming mathematical proofs and problems expressed in natural language into formal specifications and proofs in Isabelle language. The generated data was used to train a neural theorem prover, enhancing the effectiveness of the original prover. Following it, Cunningham et al. (2023) utilized an encoder-decoder framework based on the universal transformer architecture, converting both problem statements and mathematical proofs written in LaTeX into the language of the Coq interactive prover.  Jiang et al. (2022c) built a pipeline of Draft, Sketch, Prove (DSP), where the informal and incomplete proof is first generated (Draft) and given to LLMs for autoformalization (Sketch), and finally passed to off-the-shelf prover to be completed (Prove).
@@
3.1.5. Summary
When examining the progress of large language models (LLMs) in terms of their mathematical abilities, it becomes evident that the primary obstacle lies in the inherent conflict between the principles of mathematical logic and those of text generation. This discrepancy manifests itself not only in the outcomes (e.g., LLMs encountering difficulties with multiplication involving large numbers and struggling with complex mathematical problems) but also in the training data itself. Mathematical problems, in their symbolic form, represent only a minor fraction of the training corpora used for these expansive models. Consequently, the current approaches to bolstering the mathematical capabilities of LLMs can be broadly categorized into two main strategies: 1) Data enhancement: The most straightforward method to improve LLMs’ performance on mathematical tasks is to provide them with high-quality, relevant data during the fine-tuning phase of their training. By exposing the models to a more comprehensive and representative set of mathematical problems, their ability to handle such challenges can be significantly enhanced. 2) Tool integration: Another effective approach is to leverage external tools, such as calculators and code compilers, to compensate for the inherent limitations of LLMs. By strategically invoking these tools at the points where the models struggle, their functional shortcomings can be effectively mitigated, allowing for a more comprehensive and accurate handling of mathematical problems.
3.2. Writing
@@ 3.2.1. Text Summarization
Text summarization is a task that requires LLMs to compress lengthy texts into concise summaries while maintaining the essential information. This process presents a significant challenge for LLMs, as they must effectively comprehend and distill the key points from a wide range of diverse content, such as news articles and texts written in multiple languages. In the context of education, students are often confronted with an overwhelming volume of intricate learning materials. A well-crafted summary can prove invaluable in helping them grasp the core concepts quickly and efficiently, saving them considerable time and effort. For instance, a summary can break down a complex piece of code into its fundamental components, making it easier for students to understand its structure and functionality. Similarly, a summary can highlight the main ideas and key takeaways from a lengthy chapter, allowing students to focus on the most critical information without getting lost in the details. It is evident that traditional fine-tuning methods are less effective with the advent of advanced LLMs (Pu et al., 2023; Liu et al., 2022b). Pu et al. (2023) and Liang et al. (2022) showed LLMs like ChatGPT initially lag behind fine-tuned models like T5 (Raffel et al., 2020) and BART (Lewis et al., 2019) in ROUGE scores (Lin, 2004) for text summarization. However, when human judges evaluate overall quality, LLMs outperform fine-tuned models and even standard human summaries, superior in aspects like factual consistency, fluency, and diversity. This discovery underscored the limitations of traditional evaluation methods and suggested a need for new paradigms to guide summarization tasks in the LLM era. For example, BRIO (Liu et al., 2022b) implemented a ranking task to foster more diverse summarizations. Furthermore, Liu et al. (2023c) utilized a GPT model based on BRIO to directly generate training data to guide the learning process of other models, which is similar to the process of RLHF (Stiennon et al., 2020).
Given the outstanding performance of LLMs in the domain of text summarization, researchers have already begun to tackle more challenging tasks. Liu et al. (2023b) benchmarked LLMs on instruction controllable text summarization. In this task, the input provided to the model consists of two components: the source article that needs to be summarized and a set of instructions in natural language that specify the desired characteristics of the summary output. The goal is to evaluate how well LLMs can generate summaries that adhere to these specific requirements. In a related study,  Shen et al. (2023) investigated whether LLMs could potentially replace human evaluators in assessing the quality of abstractive summarization. Abstractive summarization involves generating a summary that captures the main ideas of the source text while potentially using different words and phrases. The researchers found that, at present, LLMs are not capable of serving as reliable substitutes for human evaluators in this task. LLM evaluators rate each candidate system inconsistently and are dimension-dependent. Moreover, LLMs face challenges when comparing candidate summaries that have similar levels of performance. They find it difficult to make fine-grained distinctions between summaries of comparable quality, which limits their ability to provide accurate comparative assessments. The correlation between the ratings provided by LLMs and those given by human evaluators becomes lower when dealing with higher-quality summaries. Although LLMs can surpass humans in text summarization tasks, they are not without flaws. Current LLMs make fewer silly mistakes (e.g., entity confusion, irrelevant information generation) but more sophisticated ones (Pu et al., 2023). For example, they fill in the details related to but not directly supported by the source text, which is a kind of “hallucination”.  Liu et al. (2022a) tried to employ human feedback to enhance the summarization factual consistency. The dataset DeFacto they built contained human demonstrations and informational natural language feedback consisting of corrective instructions, edited summaries, and explanations with respect to the factual consistency of the summary.  Feng et al. (2023a) tried to resolve this hallucination problem by disentangling the comprehension and embellishment abilities of LLMs. It trained the embellishment to be consistent with the facts presented in the original text.
Overall, LLMs perform well in text summarization tasks, even surpassing humans in simple summaries, but this does not mean they are flawless. In the educational domain, helping students summarize learning materials should ensure there is no conflict between the summary and the original content. LLMs still face issues with hallucinations in this aspect. While these hallucination issues can be mitigated through post-processing techniques, hallucinations remain a fundamental problem with LLMs that extends beyond the task of text summarization. Addressing the issue of hallucinations in LLMs is an ongoing research challenge that requires further investigation and development of novel approaches. Until a satisfactory solution is found, it is important to exercise caution when using LLM-generated summaries in educational contexts and to have mechanisms in place to verify the accuracy and consistency of the summaries with the original learning materials.
3.2.2. Grammatical Error Correction
@@
We are well aware of the remarkable capability of LLMs to generate fluent and coherent conversations. However, from an educational perspective, the importance of producing grammatically correct dialogues cannot be overstated, especially for students learning a new language. The correctness of grammar in conversations plays a vital role in language acquisition, providing students with reliable examples to emulate and learn from. Numerous studies have evaluated the effectiveness of LLMs in grammatical error correction (GEC). Several works (Wu et al., 2023; Fang et al., 2023; Mohammed et al., 2023) first evaluated the error correction performance of closed-source LLMs such as ChatGPT. Although there exists a pronounced gap between ChatGPT and the previous state-of-the-art (Omelianchuk et al., 2020; Grammarly, 2023) models on the overall F0.5 metric, closer analysis shows that ChatGPT underperforms other models in terms of precision but far exceeds other models in terms of recall. That said, LLMs like ChatGPT are good at error detection. A detailed manual analysis of ChatGPT’s outputs revealed that, in most cases, it maintained grammar accuracy better than the previous methods. However, it often overcorrects sentences to increase diversity and fluency, resulting in a decrease in the recall score. As a result of this characteristic, LLMs perform better when evaluated on higher-order metrics such as fluency, which assess modifications to text. However, for issues requiring minimal edit corrections, they may not necessarily outperform traditional models. Some efforts have attempted to mitigate the problem of over-correction in large models by employing instruction-tuning techniques, encouraging them to make only the necessary, such as CoEdit (Raheja et al., 2023), which covered multiple text editing tasks (including GEC) by fine-tuning LLMs to integrate the capabilities brought by these tasks. GrammarGPT (Fan et al., 2023) collected grammatically incorrect sentences and performed instruction tuning on LLMs to improve the ability of positioning grammar errors.
Overall, LLMs perform well in the area of GEC, with their main issue being over-correction. In scenarios such as copywriting or article writing, this problem is not critical, as LLMs can assist individuals in correcting grammatical errors while crafting more fluent sentences. However, in educational settings, LLMs’ GEC capabilities are more often utilized to aid students in learning grammar. This requires LLMs to accurately identify grammatical errors in sentences. The issue of over-correction could potentially mislead students, making further adjustments necessary.
3.2.3. Summary
@@ 3.3. Programming
Programming is a process of writing code and correcting code if unexpected results are obtained. Incorporating LLMs in programming education is reshaping the future of AI-assisted programming learning. LLMs could play multiple roles: as instructors providing guidance, as teaching assistants offering personalized tutorials, and as collaborative coding partners. Studies like (Ma et al., 2023b) demonstrated improved performance (17%) and efficiency (13%) among programming novices using LLM-based assistants. Research from (Phung et al., 2023) focused on programming education tasks and benchmarks like (Fu et al., 2023) and (Ding et al., 2023), which were used to evaluate the effectiveness of LLMs. This section mainly discusses the LLMs’ coding capability development from two perspectives: code writing and code refinement, corresponding to the two stages in programming.
3.3.1. Code Writing
@@
Unlike natural language tasks, generating code requires a more rigorous token syntax and places higher demands on the training stage. A common method to improve LLMs’ performance in generating code is to train or fine-tune them on extensive code datasets (Chen et al., 2021a; Nijkamp et al., 2023). WizardCoder (Luo et al., 2023b) introduced the Evol-Instruct (Xu et al., 2023) method to generate complex and diverse instruction datasets of code-related tasks. To emulate the iterative process of humans repeatedly modifying and reviewing code, InCoder (Fried et al., 2023) utilized bidirectional encoding instead of left-to-right encoding. In addition to next-token prediction, training or fine-tuning code-aimed LLMs on additional code-related tasks could enhance their programming capabilities. LLMs first learn language patterns and representations from a large amount of text data through unsupervised learning. Then, they could be fine-tuned on labeled code tasks, allowing them to learn targeted code representations and gain a deep understanding of code structure and semantics based on the provided labels. CodeT5+ (Wang et al., 2023b) introduced the concepts of unimodal and bimodal alignment, increasing the model’s adaptability to function in different modes for various downstream tasks. During the bimodal alignment phase, the model synchronizes the representations of text-code pairs using multiple tasks, which improves its ability to understand and generate content across different modalities. CodeLlama (Rozière et al., 2023) also applied the multi-task objectives, including autoregression and causal infilling prediction, which achieves better performance among open models. MFTCoder (Liu et al., 2023a) utilized the Multi-Task Learning (MTL) technique and incorporated a training loss computation algorithm to alleviate the instability and imbalance of multi-task training. Given that code text has its unique syntax and structure compared to natural language text, the methods mentioned above all attempt to construct datasets of code and perform fine-tuning. This is the most direct and effective approach to enhancing the code capabilities of LLMs.
By fine-tuning on code datasets, LLMs can enhance the probability of generating a correct piece of code. Considering the human programming process, besides relying on the programmer’s coding skills, it also involves various stages such as consulting documentation, designing code frameworks, implementing and testing submodules, which entail numerous decision-making processes. Therefore, many works perceive LLMs as agents, viewing the coding process as a continuous series of decisions and external tool invocations.  Zhang et al. (2023a) attempts to enhance the effectiveness of model code generation by employing a tree search approach, without altering the parameters of LLM itself. Specifically, each token in the code is regarded as an action, with the generated code serving as the state. The LLM makes decisions step by step, while utilizing Monte Carlo Tree Search (MCTS) to calculate the value of each action (token) in the current state, thereby selecting the optimal action and significantly improving its pass rate in code generation. Similarly,  Zhou et al. (2023b) also treats LLM as an agent, where each action involves generating a complete piece of code. It employs MCTS to estimate the value of each action as well.  Shinn et al. (2024) introduced Reflexion framework, which enhances LLM agents through linguistic feedback. The method assigned LLMs the roles of both generating code as actors and evaluating it the code as evaluators. Additionally, it utilizes self-reflection to generate verbal reinforcement cues aimed at assisting the actor in self-improvement.  (Zhou et al., 2022a) introduces a document retriever as a precursor to the code generator, extracting relevant function descriptions from documents to provide external information to the LLM. This enables the generated code to utilize the latest library functions. By considering the LLM as an agent and utilizing external documents or tree search algorithms, the accuracy of the agent’s decisions can be improved without the need to update the model’s parameters, which reduces training costs. However, this approach also has a downside: it increases the time required for decision-making during the code generation process, resulting in lower inference efficiency compared to using the LLM alone.
In addition to single-agent approaches, multi-agent systems have also made significant progress in code generation tasks.  Qian et al. (2023) developed ChatDev that divided the process of writing code into four stages: designing, coding, testing, and documenting. Each stage is managed by a group of “software agents”, with the entire chat chain acting as a facilitator, assigning specific sub-tasks to each stage. The system achieved efficient code writing. Furthermore,  Hong et al. (2023) proposed MetaGPT, a Multi-agent system for coding, which assigns diverse roles to different agents, such as Product Managers, Architects, Engineers, etc. They introduce Standardized Operating Procedures (SOPs) into prompt sequences, effectively enhancing the effectiveness of code generation. Although chat-based multi-agent systems have shown significant effectiveness in code generation tasks, they impose high demands on the fundamental capabilities of LLMs due to the need for dialogue-based coordination between agents. The effectiveness of these systems generally increases with the rise in the parameter size and capability of LLMs.
3.3.2. Code Refinement
In most cases, LLMs could not generate the correct code at once. We could enable LLMs to generate a code sketch (either actual code or pseudocode) and utilize various methods to guide the model to modify and refine the code. By leveraging the inherent code correction ability of LLMs, the overall precision and quality of the code could be significantly enhanced.